OpenStudy (cenaida):

Find the slope of the line that contains the following pair of points. (3, 4), (3, -7) A. -11/6 B. -6/11 C. 0 D. Undefined

Nnesha (nnesha):

\[\huge\rm \frac{ y_2 - y_1 }{ x_2 - x_1 }\] formula to find slope use that two order pair where \[\huge\rm (x_1 , y_1 ) ( x_2 ,y_2)\] replace x and y values and then solve

OpenStudy (cenaida):

is it C?

Nnesha (nnesha):

just remember if there is 0 at the den. then undefined

OpenStudy (anonymous):

I think C

Nnesha (nnesha):

you can't divide by 0

OpenStudy (cenaida):

then its undefined?

Nnesha (nnesha):

\[\frac{ 0 }{ 1} = 0~~~~~~~~~~~~` \frac{ 1 }{ 0 } = undfined \]

OpenStudy (anonymous):

Amen brother

OpenStudy (cenaida):

then any fraction that has a 0 will be undefined?

Nnesha (nnesha):

did you get 0 at the den. ??

OpenStudy (cenaida):

i got 0/-11

Nnesha (nnesha):

if u get 0 at the "denominator" then yes undefined

OpenStudy (cenaida):

but its on the top

Nnesha (nnesha):

how did you replaced x and y values ??

Nnesha (nnesha):

show your work please :)

OpenStudy (anonymous):

the first digit before the coma in a pair like (1,2) is x.

OpenStudy (anonymous):

so it is in (x,y) format

OpenStudy (cenaida):